# Exploring Influence Factors on LLM Suitability for No-Code Development of End User IoT Applications

### v1

@@ Abstract.
With the increasing popularity of IoT applications, end users demand more personalized and intuitive functionality. A major obstacle for this, however, is that custom IoT functionality today still requires at least some coding skills. To address this, no-code development platforms have been proposed as a solution for empowering non-technical users to create applications. However, such platforms still require a certain level of technical expertise for structuring process steps or defining event-action relations. The advent of LLMs can further enhance no-code platforms by enabling natural language-based interaction, automating of complex tasks, and dynamic code generation. By allowing users to describe their requirements in natural language, LLMs can significantly streamline no-code development. As LLMs vary in performance, architecture, training data used, and the use cases they target, it is still unclear which models are best suited and what are the influence factors determining this fit. In particular, no-code development of IoT applications by non-technical users will have completely different demands on LLMs than, e.g., code generation for more open-ended applications or for supporting professional developers. In this paper, we explore the factors influencing the suitability of LLMs to no-code development of IoT applications. We also examine the role of input prompt language on accuracy and quality of generated applications as well as the influence of LLM training data. By conducting comprehensive experiments with a range of LLMs, we provide valuable insights for optimizing LLM-powered no-code platforms, guiding the selection of the suitable LLMs and their effective application. Our findings contribute to improving the accessibility, efficiency, and user experience of no-code IoT development, ultimately enabling broader adoption of IoT technologies among non-expert users.
Large Language Models, No-Code Development Platforms
@@ 1. Introduction
The Internet of Things (IoT) is transforming how people interact with the world by seamlessly integrating smart devices into daily life. For example, smart home applications enable users to remotely control and automate household devices, while smart healthcare applications help individuals monitor their health and fitness through wearable devices (Werner et al., 2017; Pallas et al., 2020). Personalizing IoT applications enhances usability and effectiveness by delivering smart experiences tailored to users’ unique needs and lifestyles. In smart home settings customizations include, e.g., the configurations of device behavior, automation rules and third-party integrations. For example, some users may prefer window blinds that automatically adjust based on sunlight exposure to maintain steady indoor lighting, while others might prioritize wake-up routines synchronized with natural light or security systems that adapt to their daily schedules. These personalized configurations not only enhance convenience and comfort but can also improve energy efficiency, security, and privacy (Werner et al., 2017; Bao et al., 2016).
@@
Different end users have unique preferences, routines, and priorities, which can be difficult to capture in a one-size-fits-all IoT application. Achieving tailored functionality often requires coding knowledge, making it inaccessible to non-technical users. No-code development platforms have emerged as a solution, enabling end users to create custom applications without requiring any technical expertise (Blackstock and Lea, 2016). Although no-code platforms significantly simplify the application development process, users still need to navigate complex interfaces and workflows, understand platform-specific components and abstractions, e.g., (Bohé et al., 2021; Dantas et al., 2019; Younes et al., 2018; Batista et al., 2018), which makes widespread adoption difficult, especially for those unfamiliar with structured automation concepts. While rule-based platforms, e.g., IFTTT, reduce this barrier by providing basic automation, true interoperability across different ecosystems typically depends on API integration and programming skills. This reliance on technical skills further limits accessibility to non-technical users.
@@
For this limitation, today’s Large Language Models (LLMs) present new opportunities for lowering the technical knowledge requirement for IoT application development (Wang et al., 2025). By leveraging natural language processing and code generation capabilities, LLMs can transform user intentions into functional codes (Wang et al., 2025), i.e., a natural language description of desired functionality is the only information required from the end users. Different LLMs vary regarding their architecture, capability, training data, and design focus. While leaderboards [^1] For example: https://artificialanalysis.ai/leaderboards/models, https://llm-stats.com/, http://huggingface.co/open-llm-leaderboard, https://huggingface.co/spaces/lmarena-ai/chatbot-arena-leaderboard compare models on general purpose tests regarding, e.g., knowledge, reasoning, or code generation, a more detailed comparison is necessary to understand the influence of varying model factors for specific tasks. For instance, no-code development of IoT applications by non-technical end users will have completely different demands on LLMs than code generation as part of a professional developer’s toolchain. To optimize the development process of no-code IoT platforms, it is critical to understand model performance differences for the task of translating natural language user input into code for IoT customization.
Specifically, in the end user IoT domain, leveraging flexible, fine-grained, and high-level abstractions from the Function-as-a-Service (FaaS) paradigm for handling code execution and operation, enables LLMs to solely focus on core functionality generation (Wang et al., 2025). Existing research has so far primarily focused on the general feasibility of integrating LLMs into no-code environments (Cai et al., 2023; Chen et al., 2024; Hagel et al., 2024; Gao et al., 2024; Wang et al., 2025), but a systematic analysis of the factors influencing the performance and suitability for different tasks is still missing.
In this paper, we systematically analyze the impact factors affecting the performance of LLM-powered no-code development for end users’ custom IoT applications. We take smart home automation as a representative IoT application domain for this and investigate the following:
@@
    The impact of the LLM choice on performance considering varying task complexity, i.e., whether certain models are better suited for IoT-specific development tasks and user interaction.
@@
    The robustness of expectable no-code platform performance for different languages, namely, Chinese and English.
@@
    The role of LLM linguistic backgrounds which the text corpora a model trained on, in shaping usability and effectiveness of the no-code development platform.
By evaluating these factors, we aim to provide insights into the design and optimization of future LLM-powered no-code IoT platforms, making IoT development more accessible and enabling non-technical users to fully harness the potential of IoT applications.
The rest of the paper is organized as follows: Section 2 gives an overview of the concepts of LLMs, no-code development platforms, and FaaS for IoT applications; it also presents related work. Section 3 describes the methodology and study design we followed while investigating the impact factors of LLM-powered no-code IoT application development and Section 4 presents the experiment results and findings. Finally, we critically discuss our findings in Section 5 before coming to a conclusion in Section 6.
@@ 2. Background and Related Work
LLMs and no-code development platforms are revolutionizing the development and deployment of IoT applications by simplifying the design process and enhancing user accessibility, while Function-as-a-Service (FaaS) further streamlines automation by enabling scalable, event-driven execution of IoT tasks.
@@ 2.1. Large Language Models
LLMs are advanced artificial intelligence models trained on vast amounts of text data to understand, process and generate human-like responses, e.g., GPT serises from OpenAI (Achiam et al., 2023), Gemini from Google (Team et al., 2023), and Llama from Meta (Dubey et al., 2024). They use deep learning techniques, particularly transformer architectures, to analyze and predict text patterns, making them highly effective for various natural language processing tasks such as translation, summarization, and providing contextually relevant responses. This makes LLMs powerful tools for understanding user’s natural language input and generating corresponding outputs for various tasks, including automation workflows in IoT applications (Ulfsnes et al., 2024; Kök et al., 2024). Their ability to translate natural language into structured logic and automate tasks makes LLMs particularly beneficial for no-code development platforms, (Chen et al., 2024; Hagel et al., 2024) enabling non-technical users to interact with systems through conversational prompts. Essentially, integrating LLMs into the no-code development platforms, the barrier to customized software development is significantly lowered, as users can generate functional code, debug issues, and optimize application logic via intuitive, language-based interfaces (Cai et al., 2023; Paliwal et al., 2024; Martins et al., 2023; Gao et al., 2024).
2.2. No-Code Development Platforms for IoT Applications
@@
No-Code development platforms enable users to design, deploy, and modify applications through visual interfaces and pre-built drag-and-drop components, eliminating the need for coding expertise (Bohé et al., 2021; Dantas et al., 2019; Younes et al., 2018; Batista et al., 2018; El Kamouchi et al., 2023). The no-code platforms democratize software development by allowing individuals without programming knowledge to build functional applications efficiently. This approach streamlines the development process, leading to fast deployment, reducing costs, and enabling flexibility (Upadhyaya, 2023; Sufi, 2023). IoT applications encompass interconnected and diverse hardware and software components that generate and exchange data, adding complexity to the development process. No-code platforms offer advantages in the IoT domain by providing simplified integration, real-time monitoring, and automated workflows (Chen et al., 2022).
@@
The rapid development of LLMs has opened new avenues for enhancing no-code platforms, particularly in the context of IoT. With LLM-powered no-code IoT development platforms, non-technical users can describe their desired functionalities in natural language, LLM either generate the intermediate artifacts or provide functional codes to support the application development (Martins et al., 2023; Gao et al., 2024; Wang et al., 2025).
2.3. Function-as-a-Service(FaaS) for IoT Applications
Function-as-a-Service (FaaS) is a serverless computing model that allows developers to deploy individual functions that respond to specific events or triggers, without the need to manage servers or underlying infrastructure (Bermbach et al., 2021). FaaS is highly beneficial for IoT applications due to its event-driven nature (Pfandzelter and Bermbach, 2019), i.e., events from sensors or devices can trigger automated actions, such as data processing, sending notifications, or activating other devices.
@@ 2.3. Function-as-a-Service(FaaS) for IoT Applications
As FaaS also comes with a simple programming model, i.e., functions are typically stateless, small, and event-driven with the provider handling all operational complexity, it also significantly reduces the complexity LLMs have to handle when generating custom applications based on natural language prompts of non-technical users (Wang et al., 2025). As a result, the combination of LLMs and FaaS, thus, has the potential to disrupt the existing range of no-code platforms.
@@ 2.4. Related Work
There is research exploring the adoption of LLMs into no-code and low-code development platforms to enhance the usability, efficiency and accessibility of development.
Cai et al. (Cai et al., 2023) present an LLM-based low-code approach which incorporates multiple predefined visual programming interactions to achieve stable and controllable LLM responses. LLMs here are used to generate flowchart based on user natural language input, and users, in return, can refine the generated flowchart multiple rounds to ensure the correctness of presented workflow. Their use cases showcase the feasibility of integrating LLMs to provide solutions for text generation, solving coding tasks and providing human-like services. Gao et al. (Gao et al., 2024) propose a no-code IoT development platform by letting LLM generate trigger-action programs for IoT application, specifically, smart home application, achieving both token efficiency and result accuracy. Chen et al. (Chen et al., 2024) introduce an LLM-powered zero-code platform for template-bounded automated machine learning framework. LLM are used to enable the end-to-end model development and provide comprehensive analytical report to foster user understanding. Hagel et al. (Hagel et al., 2024) propose using LLMs to generate models from technical descriptions, improving time efficiency compared to manual modeling. The generated models are applied to the domain specific languages-based low-code platform for application customization, thus, turning the low-code platform to no-code. Their study involves technical participants working under supervision to ensure correct model generation which arises concerns about the accuracy of models generated without supervision.
@@
These low-code and no-code approaches demonstrated the feasibility of integrating LLMs into development platforms to enhance usability and efficiency. Additionally, beyond feasibility, the underlying factors that can affect system performance are not investigated in these studies yet. This remains a critical next step question in LLMs’ suitability for no-code or low-code development, i.e., what affects the performance of LLM-powered no-code and low-code platforms.
@@ 3. Methodology
We analyze the factors influencing LLM-empowered no-code development platforms through base platform selection(Section 3.1), experimental design (Section 3.2), dataset (Section 3.3), and evaluation metrics (Section 3.4). Specifically, our experiments focus on three key aspects, i.e., model selection, variation in prompt language, and community background of LLMs.
3.1. Base No-Code Platform Selection
To assess the impact of LLMs on no-code development, we select LLM4FaaS as the base platform to perform evaluation. Specifically, LLM4FaaS leverages the high levels of abstraction of FaaS paradigm to handle code execution and operation, enabling LLMs a sole focus on core functionality generation. We reuse and extend the LLM4FaaS dataset to conduct the experiments.
@@ 3.2. Experiment Plan
To systematically evaluate the impact of LLM selection used for no-code development in IoT platforms, we conduct experiments considering multiple aspects. We select five LLMs with different strong suits in, i.e., design focus, model size, domain-specific optimization, to evaluate and compare their performance in terms of syntactic and semantic success of code production. Additionally, as the user inputs can influence the understanding of LLMs, we evaluate the model performance with two different user input languages, i.e., in Chinese and English. Finally, as LLMs trained in different linguistic environments may exhibit varying performance across languages, we explore the LLMs performance by evaluating three mainstream Chinese LLMs.

### v2

# Exploring Influence Factors on LLM Suitability for No-Code Development of End User Applications

@@ Abstract.
No-Code Development Platforms (NCDPs) empower non-technical end users to build applications tailored to their specific demands without writing code. While NCDPs lower technical barriers, users still require some technical knowledge, e.g., to structure process steps or define event-action rules. Large Language Models (LLMs) offer a promising solution to further reduce technical requirements by supporting natural language interaction and dynamic code generation. By integrating LLM, NCDPs can be more accessible to non-technical users, enabling application development truly without requiring any technical expertise.
Despite growing interest in LLM-powered NCDPs, a systematic investigation into the factors influencing LLM suitability and performance remains absent. Understanding these factors is critical to effectively leveraging LLMs capabilities and maximizing their impact. In this paper, we investigate key factors influencing the effectiveness of LLMs in supporting end-user application development within NCDPs. By conducting comprehensive experiments, we evaluate the impact of four key factors, i.e., model selection, prompt language, training data background, and an error-informed few-shot setup, on the quality of generated applications. Specifically, we selected a range of LLMs based on their architecture, scale, design focus, and training data, and evaluated them across four real-world smart home automation scenarios implemented on a representative open-source LLM-powered NCDP. Our findings offer practical insights into how LLMs can be effectively integrated into NCDPs, informing both platform design and the selection of suitable LLMs for end-user application development.
Large Language Models, No-Code Development Platforms, End User Development
@@ 1. Introduction
Customizing applications enhances usability and effectiveness by delivering smart experiences tailored to users’ unique needs and lifestyles. For example, in smart home environments, some users may prefer window blinds that automatically adjust based on sunlight exposure to maintain steady indoor lighting, while others might prioritize wake-up routines synchronized with natural light or security systems that adapt to their daily schedules. These personalized configurations not only enhance convenience and comfort but can also improve energy efficiency, security, and privacy (Werner et al., 2017; Bao et al., 2016).
@@
Different end users have unique preferences, routines, and priorities, which makes it difficult to capture in a one-size-fits-all application. Achieving tailored functionality often requires coding knowledge, creating barriers for non-technical users. No-code development platforms (NCDPs) have emerged as a solution, enabling non-technical users to create custom applications without requiring any technical expertise (Blackstock and Lea, 2016). Although NCDPs significantly simplify the application development process, users still need to navigate complex interfaces and workflows, understand platform-specific components and abstractions (Bohé et al., 2021; Dantas et al., 2019; Younes et al., 2018; Batista et al., 2018), which makes widespread adoption difficult, especially for those unfamiliar with structured automation concepts. While rule-based platforms, e.g., IFTTT, reduce this barrier by providing basic automation, true interoperability across different ecosystems typically depends on API integration and programming skills. This reliance on technical skills further limits accessibility to non-technical users.
@@
For this limitation, Large Language Models (LLMs) present new opportunities for lowering the technical knowledge requirement for application development (Wang et al., 2025). By leveraging natural language processing and code generation capabilities, LLMs can transform user intentions into functional codes (Wang et al., 2025; Monteiro et al., 2025) - a natural language description of desired functionality is the only information required from the end users. However, variations in architecture, capabilities, training data, and design objectives among different LLMs can result in performance differences across tasks thus influence their suitability for applications development. While leaderboards compare models on general purpose tests regarding (Artificial Analysis, [n. d.]; LLM Stats, [n. d.]; Hugging Face, [n. d.]; Hugging Face and lmsys.org, [n. d.]), e.g., knowledge, reasoning, or code generation, a more detailed comparison is necessary to understand the influence of varying model factors for specific tasks. For instance, no-code development of applications by non-technical end users will have completely different demands on LLMs than code generation as part of a professional developer’s toolchain. To optimize NCDP development and broaden access, it is critical to understand LLM performance differences in translating natural language instructions into code for user-driven application customization tasks.
While the landscape of LLM-powered NCDPs is rapidly evolving, existing research has primarily focused on the feasibility of LLM integration (Cai et al., 2023; Chen et al., 2024; Hagel et al., 2024; Gao et al., 2024; Wang et al., 2025). Systematic, comparative studies of LLM performance within these platforms remain scarce. Rather than introducing platform-level variability, this paper focuses on isolating and systematically analyzing the factors that influence the suitability of LLMs for end-user application development on NCDPs. Since LLMs serve as the core and shared component across these LLM-powered NCDPs, studying their behavior allows us to derive broadly applicable, generalizable, and transferable insights. Our study investigates three critical aspects, i.e., LLM selection, prompt language, and original model training data. We conduct our evaluation using a representative LLM-powered NCDP (Wang et al., 2025), designed specifically for end users without technical backgrounds, and build on Function-as-a-Service (FaaS) abstractions. This architecture decouples infrastructure handling concerns from function logic generation, which has been shown to enhance the reliability of LLM outputs, making it server as a stable and representative foundation for isolating and examining LLM-specific factors in NCDPs. We evaluate eight LLMs on four smart home automation tasks based on real non-technical user description and investigate the following:
@@
    Impact of LLM choice on performance across varying task complexities, examining whether certain models are better suited for development tasks with non-technical user interactions.
@@
    The robustness of expectable NCDP performance for different languages, namely, Chinese and English.
@@
    The role of LLM linguistic backgrounds which the text corpora a model trained on, in shaping usability and effectiveness of NCDPs.
By evaluating these factors and incorporating a syntactic error-informed few-shot approach, we aim to provide insights into the design and optimization of future LLM-powered NCDPs, making application development more accessible and enabling non-technical users to fully harness the potential of end-user applications.
The rest of the paper is organized as follows: Section 2 gives an overview of the concepts of LLMs, NCDPs, and FaaS paradigm; it also presents related work. Section 3 describes the methodology and study design we followed while investigating the impact factors of LLM-powered NCDPs, and Section 4 presents the experiment results and findings. Finally, we critically discuss our findings in Section 5 before coming to a conclusion in Section 6.
@@ 2. Background and Related Work
LLMs and NCDPs are revolutionizing the development and deployment of applications by simplifying the design process and enhancing user accessibility, while FaaS further streamlines automation by enabling scalable, event-driven task execution.
@@ 2.1. Large Language Models
LLMs are advanced artificial intelligence models trained on vast amounts of text data to understand, process and generate human-like responses. A wide range of LLMs, varying in architecture, scale, training data, and deployment context, become available, from proprietary models served via APIs to open-source models optimized for local or custom deployment, e.g., GPT series from OpenAI (Achiam et al., 2023), Gemini from Google (Team et al., 2023), and LLaMA from Meta (Dubey et al., 2024), etc. These models use deep learning techniques, particularly transformer architectures, to analyze and predict text patterns, making them highly effective for various natural language processing tasks, e.g., translation, summarization, and providing contextually relevant responses. Their ability of translating natural language into structured logic and automating tasks makes LLMs particularly beneficial for NCDPs, enabling non-technical users to interact with systems through conversational prompts (Chen et al., 2024; Hagel et al., 2024; Ulfsnes et al., 2024; Kök et al., 2024).
Essentially, integrating LLMs into the NCDPs, the barrier that end users face in customized software development is significantly lowered, as users can generate functional code, debug issues, and optimize application logic via intuitive, natural language based interfaces (Cai et al., 2023; Paliwal et al., 2024; Martins et al., 2023; Gao et al., 2024).
2.2. No-Code Development Platforms
@@
NCDPs enable users to design, deploy, and modify applications through visual interfaces and pre-built drag-and-drop components, eliminating the need for coding expertise (Bohé et al., 2021; Dantas et al., 2019; Younes et al., 2018; Batista et al., 2018; El Kamouchi et al., 2023). The NCDPs democratize software development by allowing individuals without programming knowledge to build functional applications efficiently. This approach streamlines the development process, leading to fast deployment, reducing costs, and enabling flexibility (Upadhyaya, 2023; Sufi, 2023; Chen et al., 2022).
@@
The emergence of LLMs has opened new avenues for enhancing NCDPs. With LLM-powered NCDPs, non-technical users can describe their desired functionalities in natural language, LLM either generate the intermediate artifacts or provide functional codes to support the application development (Martins et al., 2023; Gao et al., 2024; Wang et al., 2025; Monteiro et al., 2025).
2.3. Function-as-a-Service(FaaS) Paradigm
@@
Function-as-a-Service (FaaS) is a serverless computing model that allows developers to deploy individual functions that respond to specific events or triggers, without the need to manage servers or underlying infrastructure (Bermbach et al., 2021; Pfandzelter and Bermbach, 2019). As FaaS also comes with a simple programming model, i.e., functions are typically stateless, small, and event-driven with the provider handling all operational complexity, integrating it with LLMs can significantly reduce the task complexity that LLMs have to handle when generating custom applications based on natural language prompts of non-technical users (Wang et al., 2025). As a result, this combination has the promise to enhance and expand the current landscape of NCDP.
@@ 2.4. Related Work
Prior work has examined the feasibility of integrating LLMs into NCDPs (Section 2.4.1) and the factors affecting the suitability of LLMs from technical perspectives (Section 2.4.2).
2.4.1. LLM Feasibility in NCDPs
Existing work investigating the feasibility of LLMs in NCDPs varies in terms of targeted application domains, employed technical approaches, and underlying design paradigms. Some work focuses on translating natural language into structured workflows or applications, for example, Cai et al. (Cai et al., 2023) present an LLM-driven low-code approach that generates editable flowcharts from user instructions for iterative refinement; Esashi et al. (Esashi et al., 2025) target FaaS workflow generation to benefit Cloud developers, using a dataset for compositional multi-tool tasks, though excluding execution; and Wang et al. (Wang et al., 2025), our prior work, use GPT-4o to generate serverless functions and deploy them via FaaS platform, thereby enabling customized application development for non-technical users. There is also work aiming to improve LLMs’ generation quality through multiagent architectures or domain-specific prompting, for example, Gao et al. (Gao et al., 2024) decompose IoT trigger-action-program creation into subtasks coordinated by specialized agents, and Koubaa et al. (Koubaa et al., 2025) integrate ontologies into prompts for Robot Operating System (ROS2) specific command generation. Some work adapts LLM-based platforms to particular domains or examining broader impacts, for example, Monteiro et al. (Monteiro et al., 2025) propose NoCodeGPT for web application development with rollback mechanisms, Hagel et al. (Hagel et al., 2024) automate the DSL models generation based on technical descriptions, Chen et al. (Chen et al., 2024) introduce a template-bounded no-code automated machine learning (AutoML) framework, and Liu et al. (Liu et al., 2024) conduct an empirical comparison of traditional low-code programming and an LLM-based approach using GPT3, by analyzing developer discussions on Stack Overflow.
@@ 2.4.1. LLM Feasibility in NCDPs
These work confirm the feasibility of embedding LLMs in NCDPs for diverse scenarios, however, they typically assess only a single model, focusing on whether an approach works. Beyond feasibility, the underlying factors that can affect system performance are not investigated in these studies yet, leaving it unclear whether the presented feasibility would generalize to other models or conditions. This remains a critical next step question in LLMs’ suitability for no-code or low-code development, i.e., what LLM-related factors affect the performance of LLM-powered no-code and low-code platforms.
2.4.2. LLM Suitability in other Domains
There exists research examining LLM suitability in diverse technical domains, for example, Li et al. (Li et al., 2025) compare GPT models with smaller fine-tuned models for detecting self-admitted technical debt, proposing a hybrid framework to balance precision and recall; Lu et al. (Lu et al., 2025) investigate distributed training strategies for large Transformer models, analyzing performance-memory trade-offs across model architectures and parallelization methods; Petrukha et al. (Petrukha et al., 2025) introduce SwiftEval, a benchmark for assessing LLMs on native Swift programming tasks, revealing performance variations across model families and sizes.
These studies show that understanding why an LLM performs well is crucial for informed adoption, however, they focus on technical workflows and users, whose interaction patterns and error tolerance differ from NCDP contexts. Designed for non-technical users, NCDPs aim to minimize required technical operations, introducing distinct usability needs and interaction constraints, underscoring the importance of examining LLM suitability specifically in NCDPs, where effectiveness may depend on factors beyond those studied in technical domains.
@@ 3. Methodology
To investigate the factors shaping LLM-powered NCDPs, we introduce our methodology from four aspects: base platform selection (Section 3.1), experimental design (Section 3.2), dataset construction (Section 3.3), and evaluation metrics (Section 3.4). Specifically, our experiments focus on four key aspects, i.e., model selection, variation in prompt language, community background of LLMs, and the impact of a runtime syntactic error informed few-shot setting.
3.1. Base Platform Selection
To assess the impact of LLMs on no-code development, we select a base platform, LLM4FaaS, to perform evaluation. Specifically, LLM4FaaS leverages the high levels of abstraction of FaaS paradigm to handle code execution and operation, enabling LLMs a sole focus on core functionality generation. As we focus on the LLM-related factors, we expect the base platform with a clear architecture so that the LLM behavior can be isolated and easily observed. Notably, we do not perform a comparative analysis of different NCDPs, and discuss this in Section 5.4.2. We reuse and extend the LLM4FaaS dataset to conduct the experiments and also discuss the dataset choice in Section 5.4.1.
@@ 3.2. Experiment Plan
To systematically evaluate the impact of LLM selection for NCDPs, we conduct experiments considering multiple aspects. We select five LLMs with different strong suits in, i.e., design focus, model size, domain-specific optimization, to evaluate and compare their performance in terms of syntactic and semantic success of code production. Additionally, as the user inputs can influence the understanding of LLMs, we evaluate the model performance with two different user input languages, i.e., in Chinese and English. Also, as LLMs trained in different linguistic environments may exhibit varying performance across languages, we explore the LLMs performance by evaluating three mainstream Chinese LLMs. Finally, we explore the impact of zero-shot and few-shot settings to LLM-based NCDPs.